OpenStudy (anonymous):

Solve 5-4/5x=-11 check please to see what I did wrong? 5-4/5x= -11 + 5 -4/5x= -6 -4/5x= -6 -4/5 x=-6/1 - 5/4 x=-30/4 x=-7 1/2

OpenStudy (anonymous):

\[\frac{5-4}{5x}=-11\]?

OpenStudy (anonymous):

at step 3 multiply by -5/4 to cancel the coefficient in the x term. then work from there

OpenStudy (anonymous):

\[\frac{1}{5x}=-11\]

OpenStudy (anonymous):

5x*-11=1

OpenStudy (anonymous):

-55x=1 x=1/-55

OpenStudy (anonymous):

Medal ?

OpenStudy (anonymous):

or: 5-(4/5x)=-11

OpenStudy (anonymous):

yes this correct problem 5-(4/5x)=-11

OpenStudy (anonymous):

\[5-\frac{4}{5x}=-11\]? you should told me early

OpenStudy (anonymous):

\[5+11=\frac{4}{5x}\]\[16=\frac{4}{5x}\]

OpenStudy (anonymous):

sorry. i was trying to work on another problem because i have less then an half an hr to turn my assignment in

OpenStudy (anonymous):

16*5x=4 80x = 4 x =4/80\[x=\frac{1}{20}\]

OpenStudy (anonymous):

okay got it I added the real number to the wrong side thanks medal coming now. hey can u help me solve another problem please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

post it

OpenStudy (anonymous):

Solve -3(x – 10) > -x + 5.

OpenStudy (anonymous):

-3x+30>-x+5 30-5>-x+3x 25 > 2x 25/2>x

OpenStudy (anonymous):

thanks for all your help. I really appreciate it

OpenStudy (anonymous):

urwc

OpenStudy (anonymous):

hey can u check this for me please 5/9(x) >= - 3/4 5/9-5/9(x) >= - 3/4 -5/9 x>= -3/4(9)-5/9(4) x>= -27/36 - 20/36 x>=-7/36 the answer is x>= -7/36